Danny invested some money at 3% simple interest and some money at 7% simple interest. The amount invested at the higher rate was $3,000 more than the amoung invested at the lower rate. If the total interest on the investments for 1 year was $810, then how much did he invest at each rate.
Answer by mananth(16949)

Amount @ 3.00% x
Amount @ 7.00% x + 3000
Total interest earned = 810

3.00% x+ 7.00%(x +3000 )= 810
3.00% x+7.00% *x+210= 810
3.00% x+0.07x +210=810
0.03x-0.07x= 600
0.1 x = 600
x= 6000 Amount @ 3.00%

Amount @7.00%= 6000 + 3000
= 9000

= 0
CHECK
6000 @ 3.00% = 180
9000 @ 7.00% = 630

Total income = 810
m.ananth@hotmail.ca